Simplify 1/4 of 2y+3y (2y-y)
Answer by tommyt3rd(5050)

1/4 of 2y+3y (2y-y) same as
(2y+3y (2y-y))/4=
(2y+3y^2)/4=
y/2 + 3/4 y^2